OpenStudy (anonymous):

What is the solution to the equation sqrt2x + 8 − 6 = 4

OpenStudy (anonymous):



OpenStudy (anonymous):

move all numbers on one side - so subtract 2 from each side

OpenStudy (anonymous):

square both sides in order to remove the sqrt sign now solve for x

OpenStudy (anonymous):

\[\eqalign{ &\sqrt{2x}+8-6=4 \\ &\sqrt{2x}+2=4 \\ &\sqrt{2x}=4-2 \\ &\sqrt{2x}=2 \\ &\sqrt{2x}^2=2^2 \\ &2x=4 \\ &x=4/2=2 }\]

OpenStudy (anonymous):

i think X= 2

OpenStudy (anonymous):

That is why its good to not just give answers - we should be explaining not providing answers

OpenStudy (anonymous):

i accidentally made a typo. the sqrt is over 2x+8 and not just 2x. i'm sorry.

OpenStudy (mertsj):

1. Add 6 to both sides 2. Square both sides 3. Subtract 8 from both sides. 4. Divide both sides by 2 5. Check your answer.

OpenStudy (anonymous):

@JuanitaM can you please help

OpenStudy (anonymous):